1)Jimmy can catch 12 fish in a day or pick 18 coconuts. Robert can catch 8 fish in a day or pick 4 coconuts.

a. Who has a comparative advantage in fish?

b. If they want to get 10 fish between them, how many coconuts can they get?


2) Let the following be a production function for a fishery where L is number of fishermen and TP is total number of fish caught. The price of a fish is $5 and the wage of a fisherman in their next best job is $40.

L​1​ 2​3​4​5​6​7​8​9​10​ 11

TP​12​ 23​ 33​42​50​57​63​68​72​75​ 77

a. How many workers will come if the fishery is common property?

b. How many workers will be hired if the fishery is private property?

c. What is the efficient number of fishermen?
Answer #1

1) a) Opportunity cost of producing fish for Jimmy = 18/12 = 1.5 coconuts.

Opportunity cost of producing fish for Robert = 4/8 = 0.5 coconuts.

Opportunity cost of producing coconuts for Jimmy = 12/18 = 0.67 fish.

Opportunity cost of producing coconuts for Robert = 8/4 = 2 fishes.

So we can see that Robert has a comparative advantage in the production of fish and Jimmy has a comparative advantage in the production of coconuts.

b) As Robert has a comparative advantage in the production of fishes, Robert will catch 8 fishes. Now, we know that Jimmy will also have to produce 2 fishes in a day. . For producing 2 fishes, Jimmy has to give up on 1.5*2 = 3 coconuts. So, now the number of coconuts that will be produced in total will be 15.

IF YOU FIND THIS ANSWER HELPFUL, PLEASE UP VOTE.